OpenStudy (anonymous):

A cone has a diameter of 12 in. and a height of 16 in. What is the volume of the cone? Use 3.14 as an approximation for and express your final answer in hundredths. ____in3

OpenStudy (anonymous):

I have the formula for the volume of a cone. V = 1/3 Bh

OpenStudy (anonymous):

This applies for pyramids, as well.

HanAkoSolo (jamierox4ev3r):

wow that didn't load completely, hold on

HanAkoSolo (jamierox4ev3r):

\[V=\pi r^{2}\frac{ h }{ 3 }\]

OpenStudy (anonymous):

\[V = \frac{ 1 }{ 3}\pi \times r^2h\] where h is height, r is radius. \[V = \frac{ 1 }{ 3}\times 3.14 \times (\frac{ 12 }{ 2 })^2 \times 16 = 602.88 inc^3\]

HanAkoSolo (jamierox4ev3r):

this is the formula we're looking for

HanAkoSolo (jamierox4ev3r):

oh shoot, bora beat me to it lmao. Basically, to use this formula, you take your diameter, and divide it by 2 to find the radius

HanAkoSolo (jamierox4ev3r):

and depending on how you round, you could get a different answer. I got 603.19 But the two answers aren't super far apart. @SarahStyles do you understand how to find volume of a cone?

OpenStudy (anonymous):

Yes! Thank You @Jamierox4ev3r and @BoraSemiz :)

OpenStudy (anonymous):

I got the results back... The correct answer was 602.88. But thanks for helping :)